If f(x)=5x-1, find the following f(a-2)... I have no idea of how to go about doing this, I am new to this kind of math, I would really appreciate someone's help.
Answer by stanbon(75887)

If f(x)=5x-1, find the following f(a-2)...
----------
Comment: f(a-2) tells you to substitute a-2 in place of x:
Your Problem:
f(a-2) = 5(a-2)-1
f(a-2) = 5a-10-1
f(a-2) = 5a-11
=========
Cheers,
Stan H.